OpenStudy (pigsruleyeah):

WILL FAN AND MEDAL GENEROUSLY What is the value of the expression 10 − (fraction 1 over 2)4 ⋅ 48? 2 4 5 7

OpenStudy (pigsruleyeah):

@sammixboo @Mystiz

OpenStudy (pigsruleyeah):

\[(1/2)4*49\]

OpenStudy (mystiz):

32 x 48

OpenStudy (mystiz):

i has the best math skills

OpenStudy (pigsruleyeah):

lol no

OpenStudy (pigsruleyeah):

*48 sry

OpenStudy (mystiz):

ɥʇɐɯ ʇɐ dɐɹɔ ɯɐ I

OpenStudy (pigsruleyeah):

u r crap at math

OpenStudy (pigsruleyeah):

10-(1/2)*4*48

OpenStudy (er.mohd.amir):

10-(4/2)*48=10-2*48 =10-96=-86 so i m not good in math?

OpenStudy (pigsruleyeah):

@ShadowLegendX

OpenStudy (pigsruleyeah):

@ShadowLegendX HELP PLEASE

OpenStudy (pigsruleyeah):

@Mehek14

Mehek (mehek14):

\(10-\dfrac{1}{2}4*48\) ?

OpenStudy (pigsruleyeah):

YAY

Mehek (mehek14):

Is that how it looks?

OpenStudy (pigsruleyeah):

YES

Mehek (mehek14):

\(\dfrac{1}{2}*4=?\)

OpenStudy (pigsruleyeah):

2

Mehek (mehek14):

2 * 48 = ?

OpenStudy (pigsruleyeah):

96

Mehek (mehek14):

Hold on Can you screenshot the problem?

OpenStudy (pigsruleyeah):

HOW IM ON COMPUTER

Mehek (mehek14):

Do you have a key above Backspace that says PrtScr If not, can you draw it out?

OpenStudy (pigsruleyeah):

K

OpenStudy (pigsruleyeah):



Mehek (mehek14):

Not it makes sense

Mehek (mehek14):

\(\dfrac{1}{2} ^4 = \dfrac{1}{2}*\dfrac{1}{2}*\dfrac{1}{2}*\dfrac{1}{2}=?\)

OpenStudy (pigsruleyeah):

2

Mehek (mehek14):

no

Mehek (mehek14):

\(\bf\dfrac{1}{2}*\dfrac{1}{2}*\dfrac{1}{2}*\dfrac{1}{2}=\dfrac{1x1x1x1}{2x2x2x2}=\)

OpenStudy (pigsruleyeah):

,0625

Mehek (mehek14):

0.0625 * 48 = ?

OpenStudy (pigsruleyeah):

3

Mehek (mehek14):

10 - 3 = ?

OpenStudy (pigsruleyeah):

7